# Get the Agents Drunk: Memory Perturbations in Autonomous Agent-based Recommender Systems

### v1

@@ A.1. DrunkAgent Algorithm
Table 3. Ablation Studies of DrunkAgent
Dataset Component HR@1 HR@2 HR@3 NDCG@1 NDCG@2 NDCG@3
$\mathcal{G}_{r}$ $\mathcal{S}_{t}$ $\mathcal{S}_{u}$
CDs & Vinyl $\times$ 0.0808 0.1313 0.1919 0.0808 0.1127 0.1430
$\times$ 0.0707 0.1212 0.2222 0.0707 0.1026 0.1531
0.4040 0.4141 0.4343 0.4040 0.4104 0.4205
Office Products $\times$ 0.1327 0.1429 0.1531 0.1327 0.1391 0.1442
$\times$ 0.0612 0.0714 0.1122 0.0612 0.0677 0.0881
0.2449 0.2449 0.2551 0.2449 0.2449 0.2500
$\times$ 0.0309 0.0619 0.1237 0.0309 0.0504 0.0814
0.2268 0.2268 0.2371 0.2268 0.2268 0.2320
Table 4. Success Rates to Get the Agents Drunk
Dataset #2 #5 #10 #20
CDs & Vinyl 100% 100% 100% 100%
Office Products 100% 100% 100% 100%
Musical Instruments 100% 100% 100% 100%
A.2. Dataset Statistics
@@
Table 2 illustrates the statistics of these datasets. The datasets vary in size and sparsity, which is suitable for a comprehensive evaluation. Following the work (Zhang et al., 2024b), we further sample subsets from the datasets given the expensive API calls. Specifically, considering the importance and commonness of data sparsity and cold-start in the recommendation community, we randomly sample 100 users with 800 interactions (777 items), 99 users with 693 interactions (619 items) and 98 users with 588 interactions (483 items) from the three datasets, respectively, allowing to explore more diverse and practical interaction scenarios. Moreover, the leave-one-out strategy (He et al., 2017) is used for evaluation, where the split between the training and test set is 9:1. Scaling DrunkAgent for larger datasets is left as future work.
A.3. Ablation Studies
We remove important components of each module to conduct ablation studies of DrunkAgent. First, we consider remove the surrogate module (denoted as $\mathcal{S}_{u}$). However, if we remove it, the setting will transfer from the black-box to the white-box that is impractical and beyond the scope of this paper. Moreover, due to gpt-3.5-turbo used for victim models is closed-source and only can be called via APIs, we cannot obtain logits to optimize the adversarial triggers like Llama-3 used in the surrogate model does. Hence, we maintain the surrogate model to ensure more practical settings. We then remove the greedy optimization algorithm (denoted as $\mathcal{G}_{r}$) in the generation module to observe how the performance changes. As illustrated in Table  3, the attack performance moderately degrades after removing $\mathcal{G}_{r}$, which demonstrates the effectiveness of the optimization algorithm. The lack of dramatic performance degradation is probably due to the effectiveness of our initialization. Finally, we mask the strategy module (denoted as $\mathcal{S}_{t}$). From the table, we can find that the attack performance decreases by a large margin. This may be because memory updates during agent-environment interactions are somewhat robust to non-tailored attacks (this phenomenon can also be observed in the performance of the attack baselines). Our reported results are built upon AgentCF, and similar tendencies can be found on AgentRAG and AgentSEQ.
A.4. Parameter Sensitivity
From Table  3, we can find that the strategy module contributes more than the optimization algorithm on DrunkAgent’s attack performance. As such, we increase the maximum rounds of optimization iterations per step of victim models to evaluate the attack success rates of the attack strategy, where the default value is 2 as suggested in the paper (Zhang et al., 2024b). As illustrated in Table  4, the success rates of getting the target agents drunk are maintained at 100% regardless of the number of iterations. For the other attack parameters, based on our observations, our DrunkAgent is insensitive to them.
A.5. Victim LLM-powered Agentic Models
@@
LLM-driven agents, known for their superior autonomous interaction and decision-making capabilities, are gradually being considered as next-generation RSs. Compared with some agent-powered RSs that focus solely on user-side behavior modeling using universal LLMs (Wang et al., 2024c; Huang et al., 2024), AgentCF (Zhang et al., 2024b) emphasizes the modeling of two-sided interaction relations between user agents and item agents through the idea of collaborative filtering (CF) (He et al., 2017). Both kinds of agents are equipped with memory modules, maintaining the simulated preferences and tastes of potential adopters involving their intrinsic features and acquired behavioral information via autonomous interactions and collaborative reflection optimizations at each time step. As such, the memories of the user agents and item agents can be accordingly updated, enabling the agents to better fit the real-world interaction behaviors. To conduct personalized recommendations, AgentCF introduces various well-crafted prompting strategies on the optimized memories. We name the model according to the use of the prompting strategy, with the basic one called AgentCF, the advanced retrieval-augmented one called AgentRAG and the sequential one called AgentSEQ.
The prompt implementation and parameters of collaborative agent optimizations can be found in (Zhang et al., 2024b). For the prompting strategies, we have made some improvements for high-quality recommendations, such as Chain-of-Thought and Position Bias are introduced into the prompt template $\mathcal{P}$. The strategies are described below.
@@
    (14) $\mathcal{R}_{u}=f_{\text{AgentCF}}(\mathcal{P}\oplus m_{u}\oplus m_{t}\oplus\mathcal{C}_{t}^{u}),$
    where $m_{u}$ is the short-memory of user agent $u$, $m_{t}$ is the features of the target item agent, $\mathcal{C}_{t}^{u}$ is the feature set of the candidate item agents and $\mathcal{R}_{u}$ is the ranking result.
@@
    (15) $\mathcal{R}_{u}=f_{\text{AgentRAG}}(\mathcal{P}\oplus m_{u}\oplus m_{r}\oplus m_{t}\oplus\mathcal{C}_{t}^{u}),$
    where $m_{r}$ is retrieved specialized preference from the long-term memory of user agent $u$ by taking the memories of candidate items as queries. To ensure the quality while maintaining the efficiency, we use all-MiniLM-L6-v2555https://www.sbert.net/docs/sentence_transformer/pretrained_models.html as the sentence transformer for the similarity calculation.
@@
    (16) $\mathcal{R}_{u}=f_{\text{AgentSEQ}}(\mathcal{P}\oplus m_{u}\oplus m_{t}\oplus\mathcal{M}_{t}^{u}\oplus\mathcal{C}_{t}^{u}),$
@@
We give some examples of the prompting templates of three victim models in Table  5 for reproductions.
Table 5. Examples of prompting templates of the LLM-based agentic victim models. The blue italics represent variables such as the memories of user agents and item agents optimized via agent-environment interactions. The bond fonts indicate the Chain-of-Thought and Position Bias enhancement strategies.
@@
Table 6. Examples of baselines’ adversarial descriptions of target items on real-world datasets. The red part points out the differences from the original text.
Sentence amazing !!!
sound that will elevate your music collection.
A.6. Baselines
  • Character-level Perturbations. DeepwordBug (Gao et al., 2018) and PuncAttack (Formento et al., 2023) manipulate texts by introducing typos and inserting punctuation, respectively.
  • Word-level Perturbations. TextFooler (Jin et al., 2020) and BertAttack (Li et al., 2020), respectively, aim to replace words with synonyms or contextually similar words.
  • Sentence-level Perturbations. TrivialInsertion (Zhang et al., 2024d) inserts text content with $k$ number of positive words from a pre-defined word corpus, where the insertion is conducted at the end of the text. Here, $k=2$. ChatGPTAttack (Zhang et al., 2024d) increases the stealthiness of TrivialInsertion by rewriting the text content via GPTs. GPT-3.5-turbo is used and the maximum number of words is set to 30 here.
We give some examples of adversarial descriptions of target items generated by the baselines in Table  6. Following the works (Zhang et al., 2024b; Huang et al., 2024), the item memory of the agent can be initialized by their identity information, such as titles and categories, as shown in the benign (default) descriptions of the tables. Although the adversarial description of the target item generated from DrunkAgent is different from the benign status’s (see Fig.  2 and Table  6), such modifications are feasible due to the real-world item providers is allowed to modify their product representations via APIs (Yang et al., 2024; Cohen et al., 2021; Liu and Larson, 2021), as discussed in Section 1, especially when the benign description is just a default exemplar used for experiments. Moreover, compared with the number of queries (denoted as $\mathcal{Q}$ in the table) of the baselines, DrunkAgent has the lowest attack costs due to $\mathcal{E}$=20. The implementation and parameters of the baselines follow the open-source works: StealthyAttack [^6] https://github.com/CRIPAC-DIG/RecTextAttack (Zhang et al., 2024d) and PromptBench [^7] https://github.com/microsoft/promptbench (Zhu et al., 2023b).

### v2

@@ C.1 Supplements to Experimental Setup
Table 3: Examples of baselines’ adversarial descriptions of target items on real-world datasets. The red part points out the differences from the original text.
Sentence amazing !!!
sound that will elevate your music collection.
C.1.1 Dataset Statistics
@@
Table 2 illustrates the statistics of these datasets. The datasets vary in size and sparsity, which is suitable for a comprehensive evaluation. Following the work [5], we further sample subsets from the datasets given the expensive API calls. Specifically, considering the importance and commonness of data sparsity and cold-start in the recommendation community, we randomly sample 100 users with 800 interactions including 777 items, 99 users with 693 interactions including 619 items and 98 users with 588 interactions including 483 items from the three datasets, respectively, allowing to explore more diverse and practical interaction scenarios. Moreover, the leave-one-out strategy [44] is used for evaluation, where the split between the training and test set is 9:1. Scaling DrunkAgent for larger datasets is left as future work.
C.1.2 Baseline Methods
  • Character-level Perturbations. DeepwordBug [34] and PuncAttack [35] manipulate texts by introducing typos and inserting punctuation, respectively.
  • Word-level Perturbations. TextFooler [36] and BertAttack [37], respectively, aim to replace words with synonyms or contextually similar words.
  • Sentence-level Perturbations. TrivialInsertion [13] inserts text content with $k$ number of positive words from a pre-defined word corpus, where the insertion is conducted at the end of the text. Here, $k=2$. ChatGPTAttack [13] increases the stealthiness of TrivialInsertion by rewriting the text content via GPTs, where GPT-3.5-turbo is used as the original paper recommends and the maximum number of words is set to 30 here.
We give some examples of adversarial descriptions of target items generated by the baselines in Table 3. Following the works [5, 7], the item memory of the agent can be initialized by their identity information, such as titles and categories, as shown in the benign (default) descriptions of the tables. Although the adversarial description of the target item generated from DrunkAgent is different from the benign status’s (see Fig. 2 and Table 3), such modifications are feasible due to the real-world item providers are allowed to modify their product representations via APIs [17, 48, 20], as discussed in Section 1, especially when the benign description is just a default exemplar used for experiments. Moreover, compared with the number of queries (denoted as $\mathcal{Q}$ in the table) of the baselines, DrunkAgent has the lowest attack costs due to $\mathcal{E}$=20. The implementation and parameters of the attack baselines follow the open-source works: StealthyAttack [^3] https://github.com/CRIPAC-DIG/RecTextAttack [13] and PromptBench [^4] https://github.com/microsoft/promptbench [49].
C.1.3 Victim LLM-powered Agentic Recommender Systems
@@
LLM-driven agents, known for their superior autonomous interaction and decision-making capabilities, are gradually being considered as next-generation RSs [11]. Compared with some agent-powered RSs that focus solely on user-side behavior modeling using universal LLMs [50, 7], AgentCF [5] emphasizes the modeling of two-sided interaction relations between user agents and item agents through the idea of collaborative filtering [44]. Both kinds of agents are equipped with memory modules, maintaining the simulated preferences and tastes of potential adopters involving their intrinsic features and acquired behavioral information via autonomous interactions and collaborative optimizations at each time step, as shown on the left of Fig. 1. As such, the memories of the user agents and item agents can be accordingly updated, enabling the agents to better fit the real-world interaction behaviors. To better adapt to the dynamically changing real-world environment, the agentic RSs are equipped with real-time capabilities. Whenever new descriptions are introduced at a time step, they are integrated into the evolving memories to maintain up-to-date contextual understanding. As can be seen on the right of Fig. 1, the recommendation agents conduct personalized recommendations based on the optimized memories. To drive varied architectural designs and support diverse tasks for a comprehensive evaluation of the transferability of attacks, powerful mechanisms (e.g., retrieval augmentation) and task-specific prompts (e.g., sequential recommendation) are introduced to strengthen the agentic RSs. As such, AgentRAG and AgentSEQ are developed [5].
To maintain a fair evaluation, the implementation details of how the agents can accomplish autonomous interactions and collaborative optimizations are followed the original paper suggest [5], where gpt-3.5-turbo [^5] https://platform.openai.com/docs/overview is recommended to achieve the backbone LLMs. Specifically, such backbone delivers strong natural language understanding and generation capabilities that are sufficient for most recommendation-related tasks. In addition, it is cost-efficient, making it well-suited for multi-turn interactions where cost can quickly accumulate. Moreover, it has lower latency, supporting real-time responsiveness in interactive agent scenarios. To ensure the high-quality of recommendations, we enhance the prompting strategies for recommendation agents. Specifically, we introduce the Chain-of-Thought strategy [51] to carefully craft the prompt template $\mathcal{P}$. Given the context window constraints pointed out in [5], a target item and nine non-interactive items of each normal user are randomly selected for the ranking task of the RSs, which is also suggested by the original work. The design of recommendation agents is described below.
@@
    $\mathcal{R}_{u}=f_{\text{AgentCF}}(\mathcal{P}\oplus m_{u}\oplus m_{t}\oplus\mathcal{C}_{t}^{u}),$ (8)
    where $m_{u}$ is the short-memory of user agent $u$, $m_{t}$ denotes the features of the target item agent including metadata and the memories, $\mathcal{C}_{t}^{u}$ is the feature set of the candidate item agents and $\mathcal{R}_{u}$ is the ranking result.
@@
    $\mathcal{R}_{u}=f_{\text{AgentRAG}}(\mathcal{P}\oplus m_{u}\oplus m_{r}\oplus m_{t}\oplus\mathcal{C}_{t}^{u}),$ (9)
    where $m_{r}$ is retrieved specialized preference from the long-term memory of user agent $u$ by taking the memories of candidate items as queries. To ensure the quality while maintaining the efficiency, we use all-MiniLM-L6-v2666https://www.sbert.net/docs/sentence_transformer/pretrained_models.html as the sentence transformer for the similarity calculation.
@@
    $\mathcal{R}_{u}=f_{\text{AgentSEQ}}(\mathcal{P}\oplus m_{u}\oplus m_{t}\oplus\mathcal{M}_{t}^{u}\oplus\mathcal{C}_{t}^{u}),$ (10)
@@
We give some examples of the prompting templates of three victim models in Table 4 for reproductions.
Table 4: Examples of prompting templates of the LLM-based agentic victim models. The blue italics represent variables such as the memories of user agents and item agents optimized via agent-environment interactions. The bond fonts indicate Chain-of-Thought enhancement strategy.
@@
C.1.4 Configuration of DrunkAgent
As discussed in Section 3.1, we apply the latest open-source model LLaMA 3 for building the LLM backbone of the surrogate model $f_{\text{Agent}_{\Phi}}(\cdot)$. Its collection spans different sizes, from 8B, 70B to 405B. To enhance the efficiency and ensure the quality of the attack, we use the lightweight version: meta-llama/Meta-Llama-3-8B-Instruct [^7] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct. OpenAI’s gpt-4-turbo [^8] https://platform.openai.com/docs/models/gpt-4-turbo [52] is adopted to implement $f_{\text{LLM}_{\Psi}}(\cdot)$, because it is a state-of-the-art model with powerful language understanding, reasoning and generation capabilities. We set $\mathcal{E}=20$, |~$|\widetilde{\mathcal{M}}_{c}|=10$ and $n=5$. All random seeds in the evaluation are set to 2024 and the victim LLMs’ temperature is set to 0 for reproducibility. For all the other parameters, we keep the default values.
C.1.5 Evaluation Metrics
We use two widely-used ranking metrics to evaluate the attack transferability and stealthiness: hit ratio (HR@$\mathcal{K}\uparrow$) and normalized discounted cumulative gain (NDCG@$\mathcal{K}\uparrow$) [53, 19]. HR@$\mathcal{K}$ measures whether the target/test item is appeared in the top-$\mathcal{K}$ recommendation list of the normal users, while NDCG@$\mathcal{K}$ indicates the ranking position of the target/test item in the list. As such, a high HR@$\mathcal{K}$ and a NDCG@$\mathcal{K}$ demonstrate high attack effectiveness, and vice versa. To clearly show the performance gap among different attacks, we set $\mathcal{K}$ to 1, 2, 3. We also introduce standard sentence perplexity score $\downarrow$ [13] that is a commonly-adopted metric to evaluate the attack imperceptibility. The perplexity measures a language model’s uncertainty by taking the exponential of the average negative log-likelihood of the words in a sequence, given their preceding context. Text with low perplexity is more fluent and natural, making it harder to detect as manipulated or adversarial, while text with high perplexity is more likely to be detected as unusual or abnormal.
C.2 Supplements to Performance Comparison
C.2.1 Attack Universality across Different Target Items.
In addition to cross-model transferability, we evaluate the cross-sample transferability of DrunkAgent. We incorporate the generated adversarial descriptions into totally different target items that are randomly sampled from each dataset. HR@1 (i.e., NDCG@1) of DrunkAgent across different models and target items on real-world datasets are shown in Fig. 6, which indicates its commendable transferability and triggers’ universality across different target items. This may be attributed to the fact that our triggers encapsulate diverse and general item characteristics, and, crucially, they also reveal the intended attack objectives.
Figure 6: Attack Universality across Different Target Items
Table 5: Ablation Studies of DrunkAgent
Dataset Component HR@1 HR@2 HR@3 NDCG@1 NDCG@2 NDCG@3
$\mathcal{G}_{r}$ $\mathcal{S}_{t}$ $\mathcal{S}_{u}$
CDs & Vinyl $\times$ 0.0808 0.1313 0.1919 0.0808 0.1127 0.1430
$\times$ 0.0707 0.1212 0.2222 0.0707 0.1026 0.1531
0.4040 0.4141 0.4343 0.4040 0.4104 0.4205
Office Products $\times$ 0.1327 0.1429 0.1531 0.1327 0.1391 0.1442
$\times$ 0.0612 0.0714 0.1122 0.0612 0.0677 0.0881
0.2449 0.2449 0.2551 0.2449 0.2449 0.2500
$\times$ 0.0309 0.0619 0.1237 0.0309 0.0504 0.0814
0.2268 0.2268 0.2371 0.2268 0.2268 0.2320
Table 6: Success Rates to Get the Target Agents Drunk
Dataset #2 #5 #10 #20
CDs & Vinyl 100% 100% 100% 100%
Office Products 100% 100% 100% 100%
Musical Instruments 100% 100% 100% 100%
C.2.2 Ablation Studies
We remove important components of each module to conduct ablation studies of DrunkAgent. First, we consider remove the surrogate module (denoted as $\mathcal{S}_{u}$). However, if we remove it, the setting will transfer from the black-box to the white-box that is impractical and beyond the scope of this paper. Moreover, due to gpt-3.5-turbo used for victim models is closed-source and only can be called via APIs, we cannot obtain logits to optimize the adversarial triggers like Llama-3 used in the surrogate model does. Hence, we maintain the surrogate model to ensure more practical settings. We then remove the greedy optimization algorithm (denoted as $\mathcal{G}_{r}$) in the generation module to observe how the performance changes. As illustrated in Table 5, the attack performance moderately degrades after removing $\mathcal{G}_{r}$, which demonstrates the effectiveness of the optimization algorithm. The lack of dramatic performance degradation is probably due to the effectiveness of our initialization. Finally, we mask the strategy module (denoted as $\mathcal{S}_{t}$). From the table, we can find that the attack performance decreases by a large margin. This may be because memory updates during agent-environment interactions are somewhat robust to non-tailored attacks (this phenomenon can also be observed in the performance of the attack baselines). Our reported results are built upon AgentCF, and similar tendencies can be found on AgentRAG and AgentSEQ.
C.2.3 Parameter Sensitivity
From Table 5, we can find that the strategy module contributes more than the optimization algorithm on DrunkAgent’s attack performance. As such, we increase the maximum rounds of optimization iterations per step of victim models to evaluate the attack success rates of the attack strategy, where the default value is 2 as suggested in the paper [5]. As illustrated in Table 6, the success rates of getting the target agents drunk are maintained at 100% regardless of the number of iterations. For the other attack parameters, based on our observations, our DrunkAgent is insensitive to them.
C.2.4 Attack Transferability to Victim Agentic RSs with Powerful LLM backbones
To further evaluate the transferability of DrunkAgent, we improve the LLM backbones of victim agentic RSs with more powerful gpt-4-turbo. Due to such API calls are quite expensive, we evaluate DrunkAgent on a representative real-world dataset. To ensure a sufficiently challenging evaluation, we choose the largest dataset: CD & Vinyl. Fig. 7 shows HR@1 (i.e., NDCG@1) of the benign status and DrunkAgent against the three victim models on the dataset, which demonstrates that DrunkAgent is still transferable even if the agentic RSs become powerful. We leave experiments on additional datasets and more powerful agentic RSs for future work.
Figure 7: Attack Transferability on Powerful Victim Agentic RSs.
(a) AgentCF
(b) AgentRAG
(c) AgentSEQ
Figure 8: Attack Transferability. HR@$\mathcal{K}$ and NDCG@$\mathcal{K}$ of the Benign status and DrunkAgent against various black-box victim LLM-powered agent-based RSs on the real-world Yelp dataset.
C.2.5 Attack Generalizability across Domains
To further assess the generalizability of DrunkAgent, we conduct additional experiments on the Yelp dataset. Compared to the e-commerce-focused Amazon dataset, Yelp centers around user reviews of local businesses, including restaurants, cafes, bars, salons, repair shops, and other service-oriented establishments. It places a strong emphasis on experiential content, where user reviews play a central role in shaping recommendations. This domain is inherently different from traditional product recommendation: it often involves more subjective preferences and temporal dynamics (e.g., trends in dining or local services). These characteristics make Yelp a complementary and challenging testbed for evaluating the transferability and effectiveness of our method on different domains. Due to the expensive API calls, 772 interactions are randomly sample as a subset, including 61 users and 552 items. The sparsity of the subset is 97.71%, which is more dense than the e-commerce ones. Despite these domain-specific differences, as can be seen from Fig. 8, DrunkAgent remains highly effective: it consistently promotes the target items on black-box LLM-powered recommender agents under memory corruptions, confirming the transferability of our attack and the effectiveness of memory perturbations across domains. This cross-domain success highlights the adaptability and generalizability of DrunkAgent, demonstrating that its impact is not limited to e-commerce settings but extends to real-world platforms driven by experiential content and local service discovery.
C.2.6 Attack Interpretability
We focus on why adversarial attacks exhibit both transferability and stealthiness. The transferability arises because adversarial texts are optimized to carry rich as well as general features and explicit attack objectives (e.g., ‘top-level consideration’, ‘a must-have’, ‘prime choice’) to mislead the decisions of the victim models, making them effective across models that share similar language understanding capabilities. The stealthiness is attributed to the fact that these texts are optimized to retain meaningful semantics and coherence, and mainly perturb the memories of the target item agents.
C.2.7 Computing Resourses
All experiments were run on a institution high-performance computing cluster equipped with NVIDIA H100 GPUs (each with 96GB VRAM), using CUDA version 12.4 and NVIDIA driver version 550.127.05. Each individual experimental run (from initialization to evaluation) completed in under 2 hours, with peak GPU memory usage remaining below 30GB. These resource requirements apply across all datasets, methods, and target systems evaluated in the paper.
Appendix D Limitations
While our work demonstrates the effectiveness of memory-based attacks against agentic recommender systems in controlled experimental settings, there are several limitations to be acknowledged.
Our evaluation is conducted on publicly available benchmark datasets and simulated agent environments. Although these setups capture key characteristics of real-world agentic systems, they may not fully reflect the complexity and scale of industrial deployments. In particular, factors such as heterogeneous user behaviors, evolving content ecosystems, and industrial-scale memory management pose additional challenges that are beyond the scope of this study. Future work is needed to validate the generalizability and scalability of our approach in production environments.
In addition, our study focuses on demonstrating the attack effectiveness under a black-box threat model, assuming limited access to internal model parameters or memory states. While this setting reflects practical constraints faced by external attackers, a deeper exploration of defense mechanisms and potential mitigation strategies is left for future investigation.
Despite these limitations, we believe our findings offer valuable insights into the security risks posed by memory mechanisms in agentic recommender systems, and lay a foundation for further research on robust and trustworthy recommendation agents.
Appendix E Broader Impact
As agentic recommender systems are emerging as a promising paradigm for next-generation personalization [11, 8], our work contributes positively by proactively identifying potential security vulnerabilities in these systems. While such vulnerabilities could lead to potential negative societal impacts if misused, our research serves as a foundational step toward building more robust and trustworthy agentic RSs, enabling early risk awareness and informing effective mitigation strategies, thereby helping to minimize negative societal impacts before real-world deployment. Overall, we believe the positive impact of promoting secure and responsible development in this emerging area outweighs the potential risks.